OpenStudy (anonymous):

Complete the square of the function y = −5x2 + 10x + 12. Part A: Where is the vertex located and what is the maximum or minimum value? What is the equation of the axis of symmetry? Part B: What are the zeros of the function (as exact values expressed in radical form) and the y-intercept?

OpenStudy (anonymous):

@carlyleukhardt please help

OpenStudy (diamondboy):

can u use completing the square?

OpenStudy (diamondboy):

its simply taking 10x dividing it by 2 squaring it adding and subtracting your answers and then factor

OpenStudy (anonymous):

I dont understand how to do any of it

OpenStudy (anonymous):

Ok lets start on the first step of completing the square. You have the equation. \[5x ^{2}+10x + 12\]

OpenStudy (anonymous):

5x2+10x+12 = 0

OpenStudy (anonymous):

Now subtract the number on both sides of the equation for the one that does not have an X multiplying it

OpenStudy (anonymous):

What do you get?

OpenStudy (anonymous):

5x^2-10x=-12? that doesn't sound right

OpenStudy (campbell_st):

don't complete the square like that factor out the -5 so its \[-5(x^2 - 2x + ?)\] a perfect square is in the form \[x^2 + bx + c~~~~~Where~~~ c = (\frac{b}{2})^2\] so what is half of -2 then square that value... that's what you need to add

OpenStudy (anonymous):

-1 is half of -2 then -1 squared is i right?

OpenStudy (anonymous):

Find what c equals in the equation where a=5 b=2 and c=? when you factor out negative 5 from the equation \[a(x ^{2}+bx+c) || c=\left( \frac{ b }{ 2 } \right)^{2}\]

OpenStudy (anonymous):

So plug b into the second equation

OpenStudy (anonymous):

What do you get

OpenStudy (anonymous):

Actually a=-5

OpenStudy (campbell_st):

well if you square -1 its -1 x -1 = ?

OpenStudy (anonymous):

1

OpenStudy (campbell_st):

ok... so its now \[y = -5(x^2 -2x + 1) + 12\] so now you need to add the opposite value so add 5 and you get \[y = -5(x^2 -2x +1) + 17\] now just factor the perfect square and the vertex form of the parabola is \[y = a(x - h)^2 + k\] where (h, k) is the vertex

OpenStudy (campbell_st):

in your question the value of a is -5 so factor your equation so it matches the standard form